OpenStudy (ali2x2):

How to graph y = x + 4

OpenStudy (kropot72):

Have you studied the equation of a line? y = mx + b where m is the slope and b is the y-intercept.

OpenStudy (kropot72):

@ali2x2 Are you there?

OpenStudy (kropot72):

In the equation given (y = x + 4) the slope m = 1, and the y-intercept is at (0, 4).

OpenStudy (kropot72):



OpenStudy (kropot72):

@ali2x2 Do you understand?

imqwerty (imqwerty):

This is the equation of a straight line. So to graph the line jst take some ramdom values of x and get the corresponding values of y so by doing this u'll get coordinates (x, y) take some more coordinate s like this nd plot them on the graph nd then join them all (:

OpenStudy (usukidoll):

Another method: let x = any number.... choose -2,-1,0,1,2 and solve for y for example if my x = 3 then y=3+4=7 so (3,7) is a point.

OpenStudy (ali2x2):

Back

OpenStudy (ali2x2):

Brb again

OpenStudy (ali2x2):

Thank you @UsukiDoll @imqwerty and im @kropot72 for your help :D

OpenStudy (usukidoll):

you're welcome

OpenStudy (kropot72):

You're welcome :)

OpenStudy (ali2x2):